what is added in 50 to get 25 but the answer must be positive
Answer by confusedgirl81(14)

50 + 25i^2 = 25
i^2 = -1, so you can look at it like 50-25=25 but to keep the signs positive use the i^2 root. Do not get it confused with 50-25i^2 because that equals 75